7. A client with a history of ventricular tachycardia is to receive an infusion of lidocaine hydrochloride at a rate of 20 mcg/kg/minute. The client weighs 205lbs. How many mg of lidocaine should this client receive per hour? ____________________mg/hour (Round to second decimal place)
Answer #1

205lb= 205/2.2= 93kg

Drug= 20MCG/kg/minutes

= 20MCG/93/minutes

=1860mcg/minutes

MCG to mg

= 1860/1000

= 1.86mg/minute

= 1.86×60

=111.6mg/hour

Answer: 111.6 mg per hour.